OpenStudy (anonymous):

y= I 4x +1 I a. -1,3 b. -1,-3 c. 1,3 d. 1,-3

Directrix (directrix):

y= I 4x +1 I y = absolute value of ( 4x + 1) y will have to be posititve or 0. Will test option A ( -1,3) 3 = I 4(-1) +1 I ??? 3 = I 4(-1) +1 I ??? 3 = I-3 I ?? 3 = 3 -------> True Answer: A ( -1,3)